OpenStudy (anonymous):

find differentation of Y = √x+ 1/4 〖sin〗^2 (2x)

OpenStudy (astrophysics):

\[y = \sqrt{x} +1/2 \sin^2(2x)?\] if so you will have to use the chain rule

OpenStudy (anonymous):

chain rule ? separated the sqrt x and the other one you mean ?

OpenStudy (astrophysics):

I don't know your question isn't very clear

OpenStudy (anonymous):

I think he or she meant it as: \[y = \sqrt{x + \frac{1}{4} \cdot \sin^2(2x)}\]

OpenStudy (anonymous):

Sorry, I mean firstly tell your exact question..

OpenStudy (anonymous):

What is the part under square root?

rvc (rvc):

quotient rule

rvc (rvc):

use quotient rule

OpenStudy (anonymous):

\[\sqrt{x} + \frac{ 1 }{ 4 } \sin ^{2} (2x)\]

OpenStudy (astrophysics):

No need for quotient rule, chain rule and power rule is all you need

OpenStudy (anonymous):

what's the \[\frac{ dy }{ dx }\] for this question ?

OpenStudy (astrophysics):

This is implicit differentiation

OpenStudy (astrophysics):

So every time you take a derivative of y, you also have y' or dy/dx

rvc (rvc):

oh so this is the question

OpenStudy (anonymous):

That is great..

OpenStudy (anonymous):

if i use chain rule , how to differentiate the \[\sin^{2}\] ?

OpenStudy (anonymous):

Firstly tell what is differentiation for \(\sqrt{x}\) ??

OpenStudy (astrophysics):

 think of it as such